# Improved FPT Approximation Algorithms for TSP

### v1

@@ 4.1.2 The sub-algorithm: CONNECT
The main idea of guessing $\mathcal{R}^{\prime}$ is as follows. It can be verified that augmenting $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{F}$ by adding the set of good chains $\mathcal{R}$ ensures a connected graph $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}\cup\mathcal{F}$. However, since $|V(\mathcal{R})|=|V_{g}|=n-q$, we may not be able to guess $\mathcal{R}$ in FPT time. To overcome this issue, we turn to guess the set of trees $\mathcal{F}_{R}\subseteq\mathcal{F}$ connected by $R$ for each $R\in\mathcal{R}$, as $|\mathcal{F}|\leq q$ by (3). Then, we obtain a complete graph $\widetilde{G}=(\widetilde{V},\widetilde{E},\widetilde{w})$ by contracting each tree $F\in\mathcal{F}$ into a vertex $u_{F}$ and setting the weight between $u_{F}$ and $u_{F^{\prime}}$ as $\widetilde{w}(u_{F},u_{F^{\prime}})=\min_{v\in V(F),v^{\prime}\in V(F^{\prime})}w(v,v^{\prime})$. We observe that for each chain $R\in\mathcal{R}$, by shortcutting the internal vertices, we can obtain a minimal chain $R^{*}$ connecting all trees in $\mathcal{F}_{R}$ with a non-increasing weight, where $\widetilde{V}_{R}=\{u_{F}\mid F\in\mathcal{F}_{R}\}$. Moreover, $R^{*}$ corresponds to a TSP path/tour in $\widetilde{G}[\widetilde{V}_{R}]$. Thus, we can use the DP method Bellman (1962); Held and Karp (1962) to compute an optimal TSP path/tour in $\widetilde{G}[\widetilde{V}_{R}]$ with a weight of at most $w(R^{*})\leq w(R)$, which corresponds to a set of edges in $G$, denoted by $E_{R}$, connecting all trees in $\mathcal{F}_{R}$ (see (e) in Fig.3). It holds that $w(E_{R})\leq w(R^{*})$. Therefore, we set $\mathcal{R}^{\prime}=\bigcup_{R\in\mathcal{R}}E_{R}$. The details are described in CONNECT.
@@
2:Obtain a complete graph $\widetilde{G}=(\widetilde{V},\widetilde{E},\widetilde{w})$ by contracting each $F\in\mathcal{F}$ into a vertex $u_{F}$, where, for any $u_{F},u_{F^{\prime}}\in\widetilde{V}$, define $\widetilde{w}(u_{F},u_{F^{\prime}})=\min_{v\in V(F),v^{\prime}\in V(F^{\prime})}w(v,v^{\prime})$
3:for each good chain $R\in\mathcal{R}$ with $V(R)\geq 2$ do
4:    Denote the pair anchors w.r.t. $R$ by $x$ and $x^{\prime}$.
5:    Guess the set of trees $\mathcal{F}_{R}\subseteq\mathcal{F}$ connected by $R$, and let $\widetilde{V}_{R}=\{u_{F}\mid F\in\mathcal{F}_{R}\}$.
6:    Using the DP Bellman (1962); Held and Karp (1962), find a minimum TSP tour in $\widetilde{G}[\widetilde{V}_{R}]$ if $F_{x}=F_{x^{\prime}}$, and a minimum TSP path between $u_{F_{x}}$ and $u_{F_{x^{\prime}}}$ in $\widetilde{G}[\widetilde{V}_{R}]$ otherwise.
7:    Denote the set of edges in $G$ w.r.t. the TSP tour/path by $E_{R}$.
8:    Update $\mathcal{R}^{\prime}\coloneqq\mathcal{R}^{\prime}\cup E_{R}$.
9:end for
@@
CONNECT can compute $\mathcal{R}^{\prime}$, a set of edges between good vertices, in $\mathcal{O}(2^{q^{2}})$ loops such that
@@ 5 Conclusion
Lemma 9. CONNECT can compute $\mathcal{R}^{\prime}$, a set of edges between good vertices, in $\mathcal{O}(2^{q^{2}})$ loops such that
@@
Since $|\mathcal{F}|\leq q$ by (3), each good chain $R\in\mathcal{R}$ connects at most $q$ trees in $\mathcal{F}$, i.e., $|\mathcal{F}_{R}|\leq q$. Moreover, since $|\mathcal{R}|\leq q$ by (2), to guess $\mathcal{F}_{R}$ correctly for each $R\in\mathcal{R}$, CONNECT needs to try $(2^{q})^{q}$ possibilities, and thus there exists $\mathcal{O}(2^{q^{2}})$ loops. Moreover, given $\mathcal{F}_{R}$ for each $R\in\mathcal{R}$, since CONNECT computes a TSP tour/path by using the DP method Bellman [1962]; Held and Karp [1962], this takes $\mathcal{O}(q^{2}\cdot 2^{q})$ time. Thus, each loop takes $\mathcal{O}(q^{2}\cdot 2^{q})\cdot n^{\mathcal{O}(1)}$ time.
@@
First, we prove that $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ forms a connected graph. Since $T^{*}=\mathcal{A}\cup\mathcal{B}\cup\mathcal{R}$, we know that $\mathcal{A}\cup\mathcal{B}\cup\mathcal{R}$ is a connected graph. By Lemma 8, it holds that for any $vx\in\mathcal{B}$, where $v$ is a bad vertex and $x$ is an anchor, the guessed anchor $x^{\prime}$ for $x$ satisfies $x^{\prime}\in V_{x}\subseteq V_{F_{x}}$, and then there exists a limb $vx^{\prime}\in\mathcal{B}^{\prime}$ connecting $v$ and $F_{x}$. Thus, $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}\cup\mathcal{F}$ is also a connected graph. For each $R\in\mathcal{R}$ with $V(R)\geq 2$, it connects all trees in $\mathcal{F}_{R}$. By CONNECT, there exists a set of edges $E_{R}$ w.r.t. $R\in\mathcal{R}$ connecting all trees in $\mathcal{F}_{R}$, and thus $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ forms a connected graph $G_{\mathcal{A}}$.
Then, to prove that $|Odd(G_{\mathcal{A}})\cap V(F)|$ is even for each $F\in\mathcal{F}$ and $Odd(G_{\mathcal{A}})\cap V_{a}=\emptyset$, it is sufficient to prove that each bad vertex in $G_{\mathcal{A}}$ is a 2-degree vertex, and $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}$ corresponds to an Eulerian graph, where for each $F\in\mathcal{F}$ all vertices in $V(F)$ are contacted into a single vertex. Since $\mathcal{R}^{\prime}$ and $\mathcal{F}$ only contain edges between good vertices, each bad vertex in $G_{\mathcal{A}}$ is a 2-degree vertex. Moreover, by CONNECT, it can be verified that $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}$ forms an Eulerian graph (see (e) in Fig. 3). Thus, these two conditions hold.
Last, for each $R\in\mathcal{R}$, it connects all trees in $\mathcal{F}_{R}$. Recall the minimal chain $R^{*}$, obtained by shortcutting the internal vertices in $R$, which satisfies $w(R^{*})\leq w(R)$. Since $R^{*}$ corresponds to a TSP tour/path in $\widetilde{G}[\widetilde{V}_{R}]$ and CONNECT uses the DP method to compute a minimum TSP tour/path, we know that $E_{R}$, the set of edges w.r.t. the TSP tour/path, has a weight of at most $w(E_{R})\leq w(R^{*})\leq w(R)$. Therefore, we have $w(\mathcal{R}^{\prime})\leq w(R)$. ∎
@@
Since $|S|\bmod 2=0$, by arbitrarily pairing all vertices in $S$, we obtain a set of paths $\mathcal{B}$, where each $v\sim v^{\prime}\in\mathcal{B}$ denotes the unique path between $v$ and $v^{\prime}$ on $T$. Assume that there exists two paths $v...xy...v^{\prime},u...xy...u^{\prime}\in\mathcal{B}$ sharing at least one edge $xy$. We can modify these two paths into the paths $v\sim u$ and $v^{\prime}\sim u^{\prime}$, and the total number of edges on these two paths is reduced by at least 1. Therefore, by repeatedly performing the above procedure, we can ensure that $\mathcal{B}$ is a set of edge-disjoint paths.
@@
Step 1 takes $\mathcal{O}(2^{q}\cdot q!)$ loops, where each loop takes $n^{\mathcal{O}(1)}$ time Bampis et al. [2024]. Step 2 takes $\mathcal{O}(n^{2}\log n)$ time Khachay and Neznakhina [2016]. By Lemma 8, Step 3 takes $\mathcal{O}(2^{q}\cdot q^{2q}\cdot(4q^{2})^{2q})$ loops, where each loop takes $n^{\mathcal{O}(1)}$ time. By Lemma 9, Step 4 takes $\mathcal{O}(2^{q^{2}})$ loops, where each loop takes $\mathcal{O}(q^{2}\cdot 2^{q})\cdot n^{\mathcal{O}(1)}$ time. Step 5 takes $\mathcal{O}(n^{3})$ time Lawler [1976]. By Lemma 12, Step 6 takes $\mathcal{O}(n^{2})$ time.

### v2

@@ 4.1.2 The sub-algorithm: CONNECT
The main idea of guessing $\mathcal{R}^{\prime}$ is as follows. It can be verified that augmenting $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{F}$ by adding the set of good chains $\mathcal{R}$ ensures a connected graph $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}\cup\mathcal{F}$. However, since $|V(\mathcal{R})|=|V_{g}|=n-q$, we may not be able to guess $\mathcal{R}$ in FPT time. To overcome this issue, we may consider guessing the set of trees $\mathcal{F}_{R}\subseteq\mathcal{F}$ connected by $R$ for each $R\in\mathcal{R}$, as $|\mathcal{F}|\leq q$ by (3). However, this approach may lead to $2^{\mathcal{O}(q^{2})}$ loops. To improve efficiency, we refine this method to achieve $2^{\mathcal{O}(q\log q)}$ loops.
First, we obtain a complete graph $\widetilde{G}=(\widetilde{V},\widetilde{E},\widetilde{w})$ by contracting each tree $F\in\mathcal{F}$ into a vertex $u_{F}$ and setting the weight between $u_{F}$ and $u_{F^{\prime}}$ as $\widetilde{w}(u_{F},u_{F^{\prime}})=\min_{v\in V(F),v^{\prime}\in V(F^{\prime})}w(v,v^{\prime})$. Note that for each chain $R\in\mathcal{R}$, by shortcutting the internal vertices, we can obtain a minimal chain $R_{m}$ connecting all trees in $\mathcal{F}_{R}$ with a non-increasing weight. Then, for each $R_{m}$, we obtain a new chain $R^{*}$ by shortcutting the internal vertices, ensuring that each vertex in $\widetilde{V}_{a}\coloneqq\{u_{F}\mid F\in\mathcal{F},V(F)\cap V_{a}\subseteq V^{2}_{g},V(F)\setminus V_{a}\neq\emptyset\}$ appears in only one of the paths in $\mathcal{R}^{*}\coloneqq\{R^{*}\mid R\in\mathcal{R},|V(R)|\geq 2\}$. It can be verified that $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{*}\cup\mathcal{F}$ is still a connected graph. Let $\mathcal{F}_{R^{*}}$ be the set of trees in $\mathcal{F}$ connected by $R^{*}$, and define $\widetilde{V}_{R^{*}}\coloneqq\{u_{F}\mid F\in\mathcal{F}_{R^{*}}\}$. Then, the collection of sets $\widetilde{\mathcal{V}}\coloneqq\{\widetilde{V}_{R^{*}}\cap\widetilde{V}_{a}\mid R^{*}\in\mathcal{R}^{*}\}$ forms a partition of $\widetilde{V}_{a}$, which can be guessed in $2^{\mathcal{O}(q\log q)}$ loops, since $\widetilde{V}_{a}$ is known based on $\mathcal{B}^{\prime}$ and $|\widetilde{V}_{a}|\leq|\mathcal{F}|\leq q$ by (3). Thus, if we guess $\widetilde{\mathcal{V}}$, we can obtain $\widetilde{V}_{R^{*}}\cap\widetilde{V}_{a}$ for each $R\in\mathcal{R}$ with $|V(R)|\geq 2$. Moreover, since $\widetilde{V}_{R^{*}}\setminus\widetilde{V}_{a}$ corresponds to the trees in $\mathcal{F}$ containing the anchors of $R^{*}$, it is also known based on $\mathcal{B}^{\prime}$. Thus, we can obtain $\widetilde{V}_{R^{*}}$ since $\widetilde{V}_{R^{*}}=\widetilde{V}_{R^{*}}\cap\widetilde{V}_{a}\cup\widetilde{V}_{R^{*}}\setminus\widetilde{V}_{a}$, and then we can use the DP method Bellman (1962); Held and Karp (1962) to compute an optimal TSP path/tour in $\widetilde{G}[\widetilde{V}_{R^{*}}]$ with a weight of at most $w(R^{*})\leq w(R)$, which corresponds to a set of edges in $G$, denoted by $E_{R}$, connecting all trees in $\mathcal{F}_{R^{*}}$ (see (e) in Fig.3). It holds that $w(E_{R})\leq w(R^{*})$. Thus, we set $\mathcal{R}^{\prime}=\bigcup_{R\in\mathcal{R}}E_{R}$, where $E_{R}=\emptyset$ if $|V(R)|=1$. The details are described in CONNECT.
@@
2:Obtain a complete graph $\widetilde{G}=(\widetilde{V},\widetilde{E},\widetilde{w})$ by contracting each $F\in\mathcal{F}$ into a vertex $u_{F}$, where, for any $u_{F},u_{F^{\prime}}\in\widetilde{V}$, define $\widetilde{w}(u_{F},u_{F^{\prime}})=\min_{v\in V(F),v^{\prime}\in V(F^{\prime})}w(v,v^{\prime})$.
3:Set $\widetilde{V}_{a}=\{u_{F}\mid F\in\mathcal{F},V(F)\cap V_{a}\subseteq V^{2}_{g},V(F)\setminus V_{a}\neq\emptyset\}$.
4:Guess the partition of $\widetilde{V}_{a}$, $\widetilde{\mathcal{V}}=\{\widetilde{V}_{R^{*}}\cap\widetilde{V}_{a}\mid R^{*}\in\mathcal{R}^{*}\}$, where $\widetilde{V}_{R^{*}}=\{u_{F}\mid F\in\mathcal{F}_{R^{*}}\}$, $\mathcal{F}_{R^{*}}$ is the set of trees in $\mathcal{F}$ connected by $R^{*}$, and $R^{*}$ is obtained by shortcutting the internal vertices of $R\in\mathcal{R}$, ensuring that each vertex in $\widetilde{V}_{a}$ appears in only one of the paths in =$\mathcal{R}^{*}=\{R^{*}\mid R\in\mathcal{R},|V(R)|\geq 2\}$.
5:for each good chain $R\in\mathcal{R}$ with $V(R)\geq 2$ do
6:    Denote the pair anchors w.r.t. $R$ by $x$ and $x^{\prime}$.
7:     Obtain $\widetilde{V}_{R^{*}}=\widetilde{V}_{R^{*}}\cap\widetilde{V}_{a}\cup\{u_{F_{x}},u_{F_{x^{\prime}}}\}$.
8:     Using the DP Bellman (1962); Held and Karp (1962), find a minimum TSP tour in $\widetilde{G}[\widetilde{V}_{R^{*}}]$ if $F_{x}=F_{x^{\prime}}$, and a minimum TSP path between $u_{F_{x}}$ and $u_{F_{x^{\prime}}}$ in $\widetilde{G}[\widetilde{V}_{R^{*}}]$ otherwise.
9:    Denote the set of edges in $G$ w.r.t. the TSP tour/path by $E_{R}$.
10:    Update $\mathcal{R}^{\prime}\coloneqq\mathcal{R}^{\prime}\cup E_{R}$.
11:end for
@@
CONNECT can compute $\mathcal{R}^{\prime}$, a set of edges between good vertices, in $2^{\mathcal{O}(q\log q)}$ loops such that
@@ 5 Conclusion
Lemma 9. CONNECT can compute $\mathcal{R}^{\prime}$, a set of edges between good vertices, in $2^{\mathcal{O}(q\log q)}$ loops such that
@@
Since $|\widetilde{V}_{a}|\leq|\mathcal{F}|\leq q$, to guess $\widetilde{\mathcal{V}}$, the partition of $\widetilde{V}_{a}$, CONNECT needs to try $2^{O(q\log q)}$ possibilities, and thus there exist $2^{O(q\log q)}$ loops. Based on $\widetilde{\mathcal{V}}$, the running time of CONNECT is dominated by computing a TSP tour/path in $\widetilde{G}[\widetilde{V}_{R^{*}}]$ by using the DP method Bellman [1962]; Held and Karp [1962] for each $R\in\mathcal{R}$ with $|V(R)|\geq 2$, which takes $\sum_{R^{*}\in\mathcal{R}^{*}}\mathcal{O}(|\widetilde{V}_{R^{*}}|^{2}\cdot 2^{|\widetilde{V}_{R^{*}}|})$ time. Since $\widetilde{\mathcal{V}}=\{\widetilde{V}_{R^{*}}\cap\widetilde{V}_{a}\mid R^{*}\in\mathcal{R}^{*}\}$ forms a partition of $\widetilde{V}_{a}$, we have $\sum_{R^{*}\in\mathcal{R}^{*}}|\widetilde{V}_{R^{*}}\cap\widetilde{V}_{a}|=|\widetilde{V}_{a}|\leq q$. By Step 7, $|\widetilde{V}_{R^{*}}|\leq|\widetilde{V}_{R^{*}}\cap\widetilde{V}_{a}|+2$, and then $\sum_{R^{*}\in\mathcal{R}^{*}}|\widetilde{V}_{R^{*}}|=\mathcal{O}(q)$. Thus, we have $\sum_{R^{*}\in\mathcal{R}^{*}}\mathcal{O}(|\widetilde{V}_{R^{*}}|^{2}\cdot 2^{|\widetilde{V}_{R^{*}}|})=\mathcal{O}(q^{2}\cdot 2^{q})$ time. Therefore, each loop takes $\mathcal{O}(q^{2}\cdot 2^{q})\cdot n^{\mathcal{O}(1)}$ time.
@@
First, we prove that $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ forms a connected graph. Since $T^{*}=\mathcal{A}\cup\mathcal{B}\cup\mathcal{R}$, we know that $\mathcal{A}\cup\mathcal{B}\cup\mathcal{R}$ is a connected graph. By Lemma 8, it holds that for any $vx\in\mathcal{B}$, where $v$ is a bad vertex and $x$ is an anchor, the guessed anchor $x^{\prime}$ for $x$ satisfies $x^{\prime}\in V_{x}\subseteq V_{F_{x}}$, and then there exists a limb $vx^{\prime}\in\mathcal{B}^{\prime}$ connecting $v$ and $F_{x}$. Thus, $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}\cup\mathcal{F}$ is also a connected graph. For each $R\in\mathcal{R}$ with $V(R)\geq 2$, it connects all trees in $\mathcal{F}_{R}$. Recall that $\mathcal{R}^{*}$ is obtained by shortcutting the internal vertices of chains in $\mathcal{R}$, ensuring that each vertex in $\widetilde{V}_{a}$ appears in only one of the paths in $\mathcal{R}^{*}$. It can be verified that $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{*}\cup\mathcal{F}$ remains a connected graph. By CONNECT, there exists a set of edges $E_{R}$ w.r.t. $R^{*}\in\mathcal{R}^{*}$ connecting all trees in $\mathcal{F}_{R^{*}}$, and thus $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}\cup\mathcal{F}$ forms a connected graph $G_{\mathcal{A}}$.
Then, to prove that $|Odd(G_{\mathcal{A}})\cap V(F)|$ is even for each $F\in\mathcal{F}$ and $Odd(G_{\mathcal{A}})\cap V_{b}=\emptyset$, it is sufficient to prove that each bad vertex in $G_{\mathcal{A}}$ is a 2-degree vertex, and $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}$ corresponds to an Eulerian graph, where for each $F\in\mathcal{F}$ all vertices in $V(F)$ are contracted into a single vertex. Since $\mathcal{R}^{\prime}$ and $\mathcal{F}$ only contain edges between good vertices, each bad vertex in $G_{\mathcal{A}}$ is a 2-degree vertex. Moreover, by CONNECT, it can be verified that $\mathcal{A}\cup\mathcal{B}^{\prime}\cup\mathcal{R}^{\prime}$ forms an Eulerian graph (see (e) in Fig. 3). Thus, these two conditions hold.
Last, for each $R\in\mathcal{R}$, it connects all trees in $\mathcal{F}_{R}$. Recall that the chain $R^{*}$, obtained by shortcutting the internal vertices in $R$, satisfies $w(R^{*})\leq w(R)$. Since $R^{*}$ corresponds to a TSP tour/path in $\widetilde{G}[\widetilde{V}_{R^{*}}]$ and CONNECT uses the DP method to compute a minimum TSP tour/path, we know that $E_{R}$, the set of edges w.r.t. the TSP tour/path, has a weight of at most $w(E_{R})\leq w(R^{*})\leq w(R)$. Therefore, we have $w(\mathcal{R}^{\prime})\leq w(R)$. ∎
@@
Since $|S|\bmod 2=0$, by arbitrarily pairing all vertices in $S$, we obtain a set of paths $\mathcal{B}$, where each $v\sim v^{\prime}\in\mathcal{B}$ denotes the unique path between $v$ and $v^{\prime}$ on $T$. Assume that there exist two paths $v...xy...v^{\prime},u...xy...u^{\prime}\in\mathcal{B}$ sharing at least one edge $xy$. We can modify these two paths into the paths $v\sim u$ and $v^{\prime}\sim u^{\prime}$, and the total number of edges on these two paths is reduced by at least 1. Therefore, by repeatedly performing the above procedure, we can ensure that $\mathcal{B}$ is a set of edge-disjoint paths.
@@
Step 1 takes $\mathcal{O}(2^{q}\cdot q!)$ loops, where each loop takes $n^{\mathcal{O}(1)}$ time Bampis et al. [2024]. Step 2 takes $\mathcal{O}(n^{2}\log n)$ time Khachay and Neznakhina [2016]. By Lemma 8, Step 3 takes $\mathcal{O}(2^{q}\cdot q^{2q}\cdot(4q^{2})^{2q})$ loops, where each loop takes $n^{\mathcal{O}(1)}$ time. By Lemma 9, Step 4 takes $2^{\mathcal{O}(q\log q)}$ loops, where each loop takes $\mathcal{O}(q^{2}\cdot 2^{q})\cdot n^{\mathcal{O}(1)}$ time. Step 5 takes $\mathcal{O}(n^{3})$ time Lawler [1976]. By Lemma 12, Step 6 takes $\mathcal{O}(n^{2})$ time.